# Energy-Conserving Data-Driven Discretizations for Maxwell’s Equations

### v2

@@ Abstract
We study data-driven construction of spatial discretizations for the one-dimensional Maxwell system. Given high-fidelity training data generated by a spectral discretization, we learn a linear convolution stencil that approximates the spatial derivative operator appearing in Maxwell’s equations. The stencil is obtained by solving a convex quadratic optimization problem, subject to linear constraints that enforce skew-adjointness of the discrete derivative. These constraints guarantee a semi-discrete energy identity for the resulting Maxwell system. We prove that our constraints characterize the class of skew-symmetric convolution operators and express the associated numerical wave speed and CFL restriction for the classical leapfrog scheme in terms of the learned stencil’s Fourier symbol. We then compare several convex solvers for the resulting quadratic program—projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point reference implemented in CVXPY—and evaluate the learned schemes in time-dependent one-dimensional Maxwell simulations using a Crank–Nicolson (CN) time discretization. Our numerical experiments show that (i) energy-constrained learned stencils achieve accuracy comparable to standard central differences while exactly preserving the discrete electromagnetic energy under CN time-stepping, and (ii) ADMM and interior-point methods produce nearly identical operators, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
@@ 1 Introduction
Finite-difference time-domain (FDTD) methods are a standard tool for solving Maxwell’s equations in computational electromagnetics [15]. Classical schemes, such as the Yee method, are derived by hand from the continuous equations, with spatial derivatives approximated using fixed central-difference stencils and stability enforced via analytical CFL conditions. Recently, there has been growing interest in data-driven discretizations, where discrete operators are learned from numerical or experimental data, e.g., via regression or neural networks. Examples include data-driven discretization of Burgers and Navier–Stokes equations and learned coarse-grid operators for PDEs [1].
@@
In this work we address this gap in a simple but concrete setting: the one-dimensional TE Maxwell system with periodic boundary conditions. Our goal is to learn a discrete spatial derivative operator
@@
Novelty and contributions.
Building on this context, our main contributions are as follows:
  • We formulate a convex quadratic optimization problem for learning convolution stencils for the 1D Maxwell system from spectral training data. The objective is a least-squares misfit between the learned discrete derivative and high-fidelity derivatives, while the constraints encode skew-symmetry of the stencil coefficients together with simple box bounds.
  • We show that these linear constraints are exactly equivalent to skew-adjointness of the corresponding convolution operator with respect to the discrete $L^{2}$ inner product, and hence to a semi-discrete electromagnetic energy identity. This yields, to our knowledge, the first learned Maxwell stencil with a guaranteed semi-discrete energy conservation law.
  • We characterize the numerical wave speed and associated CFL restriction of the learned discrete curl operator purely in terms of its Fourier symbol $\mu(\theta)$.
  • We implement projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point method (via CVXPY) to solve the constrained quadratic program, and we benchmark the resulting stencils in time-dependent 1D Maxwell simulations. The experiments compare PDE error, energy-constraint residuals, and energy evolution across stencil radii, and demonstrate that ADMM and interior-point methods produce nearly identical, highly accurate stencils, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
@@
Although our experiments focus on a simple 1D periodic model, the framework and energy constraints extend naturally to higher-dimensional settings and more complex material models.
Our results illustrate how even simple linear constraints in a data-driven discretization can encode important physics and restore structural properties such as energy conservation.
1.1 Related work
Data-driven discretizations.
There is a growing literature on learning discretizations and coarse-grid operators for PDEs directly from data. A prominent example is the work of [1], who learn finite-difference schemes for Burgers and Navier–Stokes equations by fitting local stencils to high-fidelity solutions while optimizing accuracy on a training set. Related approaches use regression or neural networks to construct reduced models or learned fluxes for time-dependent PDEs; see, e.g., [11, 9, 14, 12, 7]. In most of these methods, the learned operators are chosen solely to minimize an empirical error measure and are not required to satisfy structural constraints such as conservation, stability, or symmetries of the underlying PDE. In contrast, our focus is on learning a discrete Maxwell operator that provably satisfies a semi-discrete energy identity through an explicit skew-adjointness constraint on the convolution stencil.
Energy-preserving FDTD and skew-adjoint schemes.
Classical finite-difference time-domain (FDTD) methods for Maxwell’s equations, such as the Yee scheme and its variants, are carefully constructed so that discrete versions of divergence constraints and energy identities hold at the semi-discrete or fully discrete level [15, 13, 10, 8]. In particular, staggered spatial discretizations and symplectic or A-stable time integrators such as the Crank–Nicolson scheme are often used to obtain skew-adjoint or nearly skew-adjoint update operators, leading to strong stability and conservation properties [3, 4, 6]. There is also a large body of work on structure-preserving and geometric integrators for Hamiltonian ODE/PDE systems, which emphasizes invariants such as energy and symplectic form [5]. Our work is closest in spirit to these structure-preserving schemes, but differs in that the discrete spatial operator is learned from data rather than hand-designed: we encode skew-adjointness as a set of linear constraints on the stencil coefficients and solve a convex quadratic program to obtain an energy-conserving discretization.
This work in context.
@@ 1.1 Related work
To the best of our knowledge, the present paper is the first to combine data-driven stencil learning for Maxwell’s equations with linear constraints that (i) exactly characterize skew-symmetric convolution operators and (ii) yield a rigorous semi-discrete energy identity and a CFL condition expressed directly in terms of the learned stencil’s Fourier symbol. This provides a bridge between unconstrained data-driven discretizations and classical energy-preserving FDTD schemes, while remaining compatible with standard convex optimization toolchains (projected gradient, Nesterov acceleration, ADMM, and interior-point solvers).
@@ 3 Wave speed and CFL for the learned operator
which expresses the admissible time step solely in terms of the learned symbol (12). In this way the learned stencil can be interpreted within the familiar CFL framework for explicit FDTD schemes, although in our numerical experiments we will use an implicit time integrator.
@@ 4.2 Least-squares formulation
in a least-squares sense. For each grid index $i$ away from the boundaries we extract local patches
@@ 4.3 Energy constraint and box bounds
while the full equality-plus-box-constrained problem (19) is handled by ADMM and the interior-point solver in CVXPY. In our experiments the solutions of (20) produced by PG and NAG remain well inside the box $M,M]^{2R+1}$, so the difference between (20) and (19) is negligible in practice.
@@ 5.2 ADMM
To better handle the combination of equality and box constraints, we also implement an alternating direction method of multipliers (ADMM) scheme. We introduce a copy $z$ of $w$ and enforce the box constraint on $z$ only:
@@ 6.2 Learned stencils for $R=1$
For $R=1$ the true centered-difference stencil is precisely $w_{\mathrm{exact}}$ from (22).
@@
Table 1 shows that all solvers recover a stencil very close to the exact centered difference. In In Figures 1 and 2 we compare the space–time evolution and the final-time profile of the electric field $E(x,t)$ for each learned stencil against the reference FD solution. The space–time plots in Figure 1 show that the learned stencils reproduce essentially the same wave propagation as the exact centered difference, while the final-time profiles in Figure 2 are visually indistinguishable from the reference solution.
Figure 1: Space–time evolution of the electric field $E(x,t)$ for the reference central-difference stencil and the learned stencils with radius $R=1$. Each panel shows $E(x,t)$ over the full simulation time.
@@ 6.3 Error and energy residual versus stencil radius
To explore the effect of stencil size, we repeat the learning and time-integration procedure for radii $R=1,2,3,4$. Figure 3 shows the final-time relative $L^{2}$ error $\mathrm{err}(w)$ for each method as a function of $R$.
Figure 3 shows that for $R=1$ all four solvers achieve comparable accuracy, with errors on the order of $5\times 10^{-3}$. As the radius increases, the behavior diverges: the projected-gradient and Nesterov methods exhibit substantially larger errors, with PG reaching $\mathcal{O}(1)$ error and NAG reaching $\mathcal{O}(10^{-2})$ for $R=3$. In contrast, the ADMM and CVX solutions maintain errors in a narrow band between approximately $3\times 10^{-3}$ and $6\times 10^{-3}$ for all radii tested. This indicates that the second-order KKT solve in ADMM and the interior-point method in CVX are more effective at exploiting the additional degrees of freedom introduced by larger stencils, whereas the first-order projected methods struggle to fully optimize the objective within the available iteration budget.
@@ 8 Discussion and extensions
The experiments above illustrate that simple linear constraints can encode important structural information in data-driven discretizations of PDEs. In this setting, the skew-symmetry constraint $Cw=d$ ensures semi-discrete energy conservation and leads to stencils that behave like standard centered differences but are obtained purely from data.
Future work will learn staggered curl operators that preserve a discrete Poynting theorem in 2D/3D, integrate material models (Lorentz/Duffing) with structure-preserving polarization updates, and incorporate boundary treatments (ABC/CPML). On the optimization side, extending the comparison to higher-order stencils, non-quadratic objectives, and learned nonlinear fluxes would help clarify the regimes in which ADMM or interior-point methods offer the best tradeoff relative to accelerated first-order algorithms. Ultimately, combining learning with physics-motivated constraints promises accurate, stable reduced models for large-scale CEM while keeping the mathematical guarantees that make FDTD dependable [15, 5].

### v3

# Physics-Constrained Learning of Energy-Preserving Stencils for Maxwell’s Equations

@@ Abstract
We study data-driven construction of spatial discretizations for the one-dimensional Maxwell system. Using high-fidelity training data from a spectral discretization, we learn a linear convolution stencil that approximates the spatial derivative operator in Maxwell’s equations. We formulate a convex quadratic program for the stencil coefficients with linear constraints that enforce skew-adjointness of the discrete derivative; these constraints guarantee a semi-discrete electromagnetic energy identity and yield a CFL condition expressed directly in terms of the stencil’s Fourier symbol. We compare several convex solvers for the resulting quadratic program—projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point reference implemented in CVXPY—and evaluate the learned operators in time-dependent Maxwell simulations using a Crank–Nicolson (CN) discretization. Numerical experiments, including cases with nonstandard target operators and noisy training data, show that (i) energy-constrained learned stencils achieve accuracy comparable to standard central differences while exactly preserving the discrete electromagnetic energy under CN time-stepping, and (ii) ADMM and interior-point solvers produce nearly identical operators, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
@@ 1 Introduction
Finite-difference time-domain (FDTD) methods are a standard tool for solving Maxwell’s equations in computational electromagnetics [21]. Classical schemes, such as the Yee method, are derived by hand from the continuous equations, with spatial derivatives approximated using fixed central-difference stencils and stability enforced via analytical CFL conditions. Recently, there has been growing interest in data-driven discretizations, where discrete operators are learned from numerical or experimental data, for example via regression or neural networks. Examples include data-driven discretization of Burgers and Navier–Stokes equations and learned coarse-grid operators for PDEs [1].
@@
In this work we address this gap in a simple but concrete setting: the one-dimensional Maxwell system with periodic boundary conditions. Our goal is to learn a discrete spatial derivative operator
@@
Classical FDTD schemes such as the Yee method [21] are hand-derived so that the discrete curl operator is skew-adjoint, yielding exact semi-discrete energy conservation. High-order extensions and operator-splitting schemes similarly maintain this structure [3, 6].
Data-driven PDE schemes learn local stencils or coarse-grid operators directly from numerical data. Bar-Sinai et al. learn finite-difference discretizations for Burgers and Navier–Stokes, showing that learned operators can match or exceed classical schemes on specific training distributions [1]. Many follow-up works learn PDE operators or closure models [17, 11, 20, 18, 9], but typically do not enforce stability, skew-adjointness, or conservation laws.
Recent neural-operator surrogates for electromagnetics incorporate physics through architectural biases or penalty terms. Dynamic-convolution neural operators for FDTD-like EM propagation impose causal receptive fields and material dispersion structure [14]. Domain-decomposition neural operators couple learned subdomain solvers through interface penalties or consistency conditions [15, 8]. Other neural operator frameworks embed conservation via soft constraints or corrective layers [13, 12]. These approaches improve stability but enforce physical structure only approximately, through optimization.
Building on this context, our main contributions are as follows. First, we formulate a convex quadratic optimization problem for learning convolution stencils for the 1D Maxwell system from spectral training data, with an objective given by a least-squares misfit between the learned discrete derivative and high-fidelity derivatives, and linear constraints that encode skew-symmetry of the stencil coefficients together with simple box bounds. Second, we show that these linear constraints are exactly equivalent to skew-adjointness of the corresponding convolution operator with respect to the discrete $L^{2}$ inner product, and hence to a semi-discrete electromagnetic energy identity, yielding what is, to our knowledge, the first learned Maxwell stencil with a guaranteed semi-discrete energy conservation law. Third, we characterize the numerical wave speed and associated CFL restriction of the learned discrete curl operator purely in terms of its Fourier symbol $\mu(\theta)$. Finally, we implement projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point method (via CVXPY) to solve the constrained quadratic program, and we benchmark the resulting stencils in time-dependent 1D Maxwell simulations; the experiments compare PDE error, energy-constraint residuals, and energy evolution across stencil radii, and demonstrate that ADMM and interior-point methods produce nearly identical, highly accurate stencils, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
@@
Although our experiments focus on a simple 1D periodic model, the framework and energy constraints extend naturally to higher-dimensional settings and more complex material models. Our results illustrate how even simple linear constraints in a data-driven discretization can encode important physics and restore structural properties such as energy conservation.
@@
Our method imposes hard linear constraints that make the learned convolution stencil exactly skew-adjoint, guaranteeing a discrete electromagnetic energy identity independent of training accuracy or solver choice. This places our work closest to Bar-Sinai et al. but with strict structure preservation rather than penalty-based regularization, and complementary to physics-informed neural operators, which typically enforce conservation approximately. To the best of our knowledge, this is the first work to combine data-driven stencil learning for Maxwell’s equations with linear constraints that (i) exactly characterize skew-symmetric convolution operators and (ii) yield a rigorous semi-discrete energy identity and a CFL condition expressed directly in terms of the learned stencil’s Fourier symbol. This provides a bridge between unconstrained data-driven discretizations and classical energy-preserving FDTD schemes, while remaining compatible with standard convex optimization toolchains (projected gradient, Nesterov acceleration, ADMM, and interior-point solvers).
Beyond the homogeneous model, Sections 6.6 and 6.7 illustrate how the same constrained learning framework can recover nonstandard, effective operators and regularize noisy or mildly inconsistent training data.
@@ 3 Wave speed and CFL for the learned operator
which expresses the admissible time step solely in terms of the learned symbol (12). In this way the learned stencil can be interpreted within the familiar CFL framework for explicit FDTD schemes, although in our numerical experiments we will use an implicit time integrator.In this work we focus on CN, which is unconditionally stable for the skew-adjoint semi-discrete operator. A systematic comparison with leapfrog using the derived CFL bound is left for future work
@@ 4.2 Least-squares formulation
in a least-squares sense. Because we assume periodic boundary conditions, the stencil is applied identically at every grid point, with indices interpreted modulo $N$.
For each grid index $i$ we extract local patches
@@ 4.3 Energy constraint and box bounds
while the full equality-plus-box-constrained problem (19) is handled by ADMM and the interior-point solver in CVXPY. In our experiments the solutions of (20) produced by PG and NAG remain well inside the box $M,M]^{2R+1}$, so the difference between (20) and (19) is negligible in practice. In the presence of noisy or mildly inconsistent training data, the least-squares system can become ill-conditioned and produce stencils with very large coefficients; the Tikhonov term and box bounds mitigate this behavior, as illustrated in Section 6.7.
@@ 5.2 ADMM
To better handle the combination of equality and box constraints, we also implement an alternating direction method of multipliers (ADMM) [4], [5] scheme. We introduce a copy $z$ of $w$ and enforce the box constraint on $z$ only:
@@ 6 Numerical experiments
All experiments were implemented in Python using NumPy for linear algebra and SciPy for FFTs. The ADMM and projected-gradient/Nesterov solvers are written explicitly as described in Section 5. The interior-point reference solution is computed using CVXPY (version X.Y) with the SCS solver. Runtimes are reported for a single-threaded run on a (Apple M4, 16 GB RAM) using Python 3.11. Although absolute times depend on hardware and implementation details, the relative trends between first-order methods (PG/NAG) and the second-order ADMM/CVX approach remained consistent across several runs.
@@ 6.2 Learned stencils for $R=1$
In In Figures 1 and 2, we compare the space-time evolution and the final-time profile of the electric field $E(x,t)$ for each learned stencil against the reference FD solution for R =1. The final-time and space-time profiles in Figure 2 are visually indistinguishable from the reference solution.
Figure 1: Space-Time Evolution
@@
This experiment serves as a sanity check for the learning framework: for the smallest radius $R=1$, the energy-constrained optimization recovers a stencil that is essentially indistinguishable from the classical centered difference, and the corresponding Crank–Nicolson time integration reproduces the reference solution to within plotting accuracy. In other words, when the true operator is known and low order, the learned discretizations do not degrade either accuracy or energy behavior, while providing the same skew-adjoint structure that underpins exact discrete energy conservation. This $R=1$ case therefore provides a baseline for the higher-radius stencils and noisy-training experiments considered below.
@@ 6.3 Error and energy residual versus stencil radius
To explore the effect of stencil size, we repeat the learning and time-integration procedure for radii $R=1,2,3,4$. Figure 3 shows the final-time relative $L^{2}$ error $\mathrm{err}(w)$ for each method as a function of $R$. We initialize all optimizations at the zero stencil; even from this uninformative start, the energy-constrained learner recovers stencils close to classical centered differences. Figure 3 shows that for $R=1$ all four solvers achieve comparable accuracy, with errors on the order of $5\times 10^{-3}$. As the radius increases, the behavior diverges: the projected-gradient and Nesterov methods exhibit substantially larger errors, with PG reaching $\mathcal{O}(1)$ error and NAG reaching $\mathcal{O}(10^{-2})$ for $R=3$. In contrast, the ADMM and CVX solutions maintain errors in a narrow band between approximately $3\times 10^{-3}$ and $6\times 10^{-3}$ for all radii tested. This indicates that the second-order KKT solve in ADMM and the interior-point method in CVX are more effective at exploiting the additional degrees of freedom introduced by larger stencils, whereas the first-order projected methods struggle to fully optimize the objective within the available iteration budget.We believe the gap for larger R is primarily due to conditioning and the fact that PG/NAG only project onto the equality constraint, whereas ADMM/CVX solve the full KKT system with box constraints. Intersection projections (e.g., Dykstra) could close this gap but would blur the simplicity advantage of PG/NAG; we leave such variants to future work
@@ 6.4 Discrete energy conservation for learned Maxwell stencils
6.5 Dispersion analysis.
To complement the end-time $L^{2}$ error metrics, we carry out a Fourier dispersion analysis of the semi-discrete Maxwell system with Crank–Nicolson time stepping. For a given stencil $w$ with Fourier symbol $\lambda(\theta)$, the one-step Crank–Nicolson update for each Fourier mode has eigenvalue
$\mu(\theta)=\frac{1+\tfrac{\Delta t}{2}\,\lambda(\theta)}{1-\tfrac{\Delta t}{2}\,\lambda(\theta)}.$
(a) Amplification error $|\mu(\theta)|-1$.
(b) Phase dispersion $\arg\mu(\theta)/\theta$.
Figure 5: Fourier dispersion analysis of the semi-discrete Maxwell system with Crank–Nicolson time stepping: (a) amplification error and (b) phase dispersion for learned vs. centered-difference stencils.
We plot the amplification factor $|\mu(\theta)|$ and the normalized phase $\arg\mu(\theta)/\theta$ as functions of $\theta\in(0,\pi]$, and compare the learned energy-constrained stencils with standard centered-difference stencils of the same radius $R$; see Figs. 5(a)5(b).
Figure 5(a) shows that $|\mu(\theta)|\approx 1$ for all learned stencils and all wavenumbers, essentially indistinguishable from the centered-difference curves and the continuous reference. This confirms that the fully discrete schemes are non-dissipative and preserve the discrete electromagnetic energy in practice, consistent with the skew-adjointness constraints imposed on $D$.
The phase plots in Fig. 5(b) provide a more stringent test of accuracy. For $R=1,2,3$, the dispersion curves $\arg\mu(\theta)/\theta$ of the learned operators closely track those of the corresponding centered differences over the range of resolved wavenumbers, with only small deviations at the highest modes; in some mid-frequency bands the learned stencils are slightly closer to the continuous reference. Overall, this mode-by-mode dispersion analysis is more diagnostic than end-time $L^{2}$ errors and confirms that the hard skew-adjointness constraints produce learned stencils with physically meaningful wave-propagation properties. see, e.g., [21, 2]
6.6 Learning a nonstandard skew-adjoint operator
To mimic an effective spatial operator arising from unresolved heterogeneity or subgrid physics, we consider a “nonstandard” convolution stencil $w^{\star}\in\mathbb{R}^{2R+1}$ with radius $R=2$ that is skew-adjoint (i.e., satisfies the constraints (11)) but does not coincide with the usual fourth-order centered difference coefficients. In this setting the analytic form of the differential operator is not available, but we can still generate high-fidelity training trajectories by evolving Maxwell’s equations using $w^{\star}$.
We then (i) learn an energy-constrained stencil $w_{\mathrm{QP}}$ by solving the quadratic program (20) with ADMM, and (ii) compare it to the standard centered-difference stencil $w_{\mathrm{CD}}$ of the same radius. Numerically we obtain
$\frac{\|w_{\mathrm{QP}}-w^{\star}\|_{2}}{\|w^{\star}\|_{2}}\approx 6\times 10^{-8},\qquad\frac{\|w_{\mathrm{CD}}-w^{\star}\|_{2}}{\|w^{\star}\|_{2}}\approx 2.5\times 10^{-1},$
so the learned operator recovers the nonstandard stencil to nearly machine precision, while the standard stencil incurs a $25\%$ relative coefficient error.
Figure 6(a) compares the final-time electric field $E(x,T)$ produced by $w^{\star}$, $w_{\mathrm{CD}}$, and $w_{\mathrm{QP}}$. The centered-difference stencil exhibits a systematic phase and amplitude error, whereas the learned stencil is indistinguishable from the nonstandard operator. At the same time, all three stencils remain essentially energy-preserving under Crank–Nicolson time stepping (Figure 6(b)), demonstrating that the proposed method can capture effective, nonstandard differential operators while retaining a discrete Poynting theorem.
(a) Noisy Discrete Energy Simulation
(b) Noisy E(x,t)
Figure 6: Nonstandard vs Standard Skew-adjoint operator
6.7 Noisy training and ill-posed least squares
To emulate noisy or partially inconsistent experimental data, we add Gaussian noise to the time derivatives $dU/dt$ used in the regression problem (24). Specifically, we draw $\tilde{dU/dt}=dU/dt+\eta$ with $\eta$ i.i.d. $N(0,\sigma^{2})$ and form the corresponding least-squares system $(A,b_{\mathrm{noisy}})$. We then compute three stencils: (i) the standard centered difference $w_{\mathrm{CD}}$, (ii) an unconstrained least-squares stencil $w_{\mathrm{LS}}$ obtained by minimizing (24) with no constraints, and (iii) an energy-constrained stencil $w_{\mathrm{QP}}$ obtained by solving the quadratic program (20) with the same noisy data.
Although $w_{\mathrm{LS}}$ achieves a very small training error, it violates the skew-adjointness constraints and develops very large high-frequency coefficients. When used in Crank–Nicolson time stepping, the resulting Maxwell semi-discretization is no longer energy-stable, and the discrete energy $\mathcal{E}^{n}$ blows up (Figure 7(a)). In contrast, the energy-constrained stencil $w_{\mathrm{QP}}$ remains exactly skew-adjoint by construction and yields stable dynamics with energy drift comparable to the hand-derived centered difference stencil.
Figure 7(b) shows the final-time electric field $E(x,T)$ for all three operators. The solution obtained with $w_{\mathrm{QP}}$ is visually indistinguishable from the Central difference baseline, while the unconstrained LS stencil produces a completely corrupted field. This experiment illustrates a regime where training data are noisy or mildly inconsistent, and the energy constraints act as a form of physics-based regularization: they prevent overfitting to noise and guarantee a discrete Poynting theorem irrespective of the data quality.
(a) Noisy Discrete Energy Simulation
(b) Noisy E(x,t)
Figure 7: Noisy training and ill-posed least squares
Figure 8: Space-Time Plot of the Electric Field Under Noise Perturbation
@@ 8 Discussion and extensions
Our results show that simple linear constraints can encode key physics in data-driven discretizations: enforcing $Cw=d$ yields skew-adjoint spatial operators that exactly preserve a semi-discrete energy identity, while remaining close to standard centered differences in the homogeneous 1D setting.
To probe more realistic regimes, we also considered (i) a nonstandard target operator whose stencil does not match any classical centered-difference formula and showed that the constrained learner accurately recovers its space–time behavior, while the matching CD stencil incurs noticeable error, and (ii) a noisy regression experiment in which unconstrained least squares produces large coefficients and severe energy drift, whereas the energy- and box-constrained stencil stays well-conditioned, energy-stable, and faithful to the clean baseline $E(x,t)$.
These examples suggest that physics-motivated constraints can regularize noisy or mildly mis-specified regression problems and make it feasible to learn effective Maxwell operators in settings where analytic stencils are difficult to derive or tune (e.g., heterogeneous media or band-limited/experimental data). A natural next step is to extend the same constrained-learning philosophy to staggered curl operators on Yee grids in 2D/3D, enforcing block skew-symmetry to preserve a discrete Poynting theorem, and to couple the learned operators with simple dispersive material models and absorbing layers in more realistic FDTD setups.
Acknowledgements
The author gratefully acknowledges Profs. Xiao Fu and Nathan Gibson for valuable guidance and feedback throughout this work.